# Position Paper: An Integrated Perspective on Data, Metrics, and Methodology for Deep Time-Series Forecasting

### v3

@@ 5 Conclusion
This paper presents an integrated perspective on data, metrics, and methodology in the field of deep time-series forecasting, facilitated by the ProbTS tool. Our findings not only provide valuable insights into the strengths and limitations of existing methods but also highlight several underexplored areas that merit future research. In addition, we expect that our insights could spur more explorations on these fields. We also hope the ProbTS tool can become a useful resource for the community, expediting future advancements in time-series forecasting.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A More Details on ProbTS
A.1 Additional Related Work
@@
Method Fore. Horizon Metrics Arch. Design Dec. Scheme
Short Long Point Distr. General Customized Autoreg. Non-autoreg.
N-BEATS (Oreshkin et al., 2020)
Autoformer (Wu et al., 2021)
Informer (Zhou et al., 2021)
Pyraformer (Liu et al., 2022)
N-HiTS (Challu et al., 2023)
LTSF-Linear (Zeng et al., 2023)
PatchTST (Nie et al., 2023)
TimesNet (Wu et al., 2023)
DeepAR (Salinas et al., 2020)
GP-copula (Salinas et al., 2019)
LSTM NVP (Rasul et al., 2021b)
LSTM MAF (Rasul et al., 2021b)
Trans MAF (Rasul et al., 2021b)
TimeGrad (Rasul et al., 2021a)
CSDI (Tashiro et al., 2021)
SPD (Bilos et al., 2023)
TSDiff (Kollovieh et al., 2023)
This Paper
A.2 Time-series Forecasting Datasets
@@ A.3 Data Visualization
It’s important to note that these visualizations are not selectively chosen or ”cherry-picked”. We have depicted multiple time-series segments from various time steps, and the observed patterns remain consistent across these different instances.
@@
A.4 Quantifying Trend and Seasonality Strengths
Based on the intuition obtained from data visualization, we would like to quantify the strengths of trend and seasonality for a time-series segment with a predefined window size (corresponding to the prediction horizon length). Then we can quantify the trend and seasonality strengths at the dataset level by averaging over all time-series segments of a dataset.
To quantify the strengths of trend and seasonality for a fixed-length time-series segment, we draw upon methodologies outlined in the work of (Wang et al., 2006). In particular, we employed a time series decomposition model expressed as:
@@ A.4 Quantifying Trend and Seasonality Strengths
A.5 Quantifying Data Distribution Complexity
@@
Hence, we use the Jensen–Shannon divergence (Nielsen, 2019) to measure the similarity between the actual value distribution of a time-series segment and a Gaussian distribution fitted to the observed values. Short-term datasets used a window size of 30, while long-term datasets used a size of 336. By averaging the calculated divergence values across all time-series segments of a dataset, we obtain a dataset-level measure of non-Gaussianity. A larger divergence value indicates a larger deviation from a Gaussian distribution in the data.
A.6 Evaluation Metrics
@@
A.6.1 Metrics for Point Forecasts
@@
$\textrm{MAE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}|x_{i,t}-\hat{x}_{i,t}|,$
where $K$ is the number of variates, $L$ is the length of series, $x_{i,t}$ and $\hat{x}_{i,t}$ denotes the ground-truth value and the predicted value, respectively. For multivariate time series, we also provide the aggregated version:
@@
where $x^{\textrm{sum}}_{t}$ and $\hat{x}^{\textrm{sum}}_{t}$ are the summation across the dimension $K$ of $x_{i,t}$ and $\hat{x}_{i,t}$, respectively.
@@
$\textrm{NMAE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}\frac{|x_{i,t}-\hat{x}_{i,t}|}{|x_{i,t}|}.$
@@
$\textrm{NMAE}_{\textrm{sum}}=\frac{1}{T}\sum^{T}_{t=1}\frac{|x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t}|}{|x^{\textrm{sum}}_{t}|}.$
@@
$\textrm{MSE}=\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{L}_{t=1}(x_{i,t}-\hat{x}_{i,t})^{2}.$
@@
$\textrm{MSE}_{\textrm{sum}}=\frac{1}{T}\sum^{L}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}.$
@@
$\textrm{NRMSE}=\frac{\sqrt{\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{L}_{t=1}(x_{i,t}-\hat{x}_{i,t})^{2}}}{\frac{1}{K\times T}\sum^{K}_{i=1}\sum^{T}_{t=1}|x_{i,t}|}.$
@@
$\textrm{NRMSE}_{\textrm{sum}}=\frac{\sqrt{\frac{1}{T}\sum^{L}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}}}{\frac{1}{T}\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}|}.$
A.6.2 Metrics for Distributional Forecasts
@@
The Continuous Ranked Probability Score (CRPS) (Matheson & Winkler, 1976) quantifies the agreement between a cumulative distribution function (CDF) $F$ and an observation $x$, represented as:
@@
where $\mathds{I}{x\leq z}$ denotes the indicator function, equating to one if $x\leq z$ and zero otherwise.
@@
A.7 Implementation Details
ProbTSwas developed using PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). During training, we sampled 100 batches per epoch and limited training to 50 epochs, using the CRPS metric for checkpointing. All experiments employed the Adam optimizer and were run on single NVIDIA Tesla V100 GPUs with CUDA 11.3. To enable evaluation of distribution-level metrics, we conducted 100 samplings to calculate metrics on the test set.
Following the most commonly adopted settings (Zeng et al., 2023; Nie et al., 2023; Wu et al., 2023), in the long-term forecasting context, all of the models are following the same experimental setup with prediction length $T\in\{24,36,48,60\}$ for ILI dataset and $T\in\{96,192,336,720\}$ for other datasets. Note that the lookback window here is 96 for all the models, to ensure a fair comparison. In the short-term forecasting context, the length of the lookback window is the same as the forecasting horizons, which are 30 for Exchange-S dataset and Wikipedia-S dataset, and 24 for the rest, the same as (Salinas et al., 2019).

### v4

@@ 5 Conclusion
In this study, we introduced ProbTS, a benchmark tool tailored for evaluating essential forecasting needs, which facilitates a detailed comparison of various state-of-the-art models in the context of time-series forecasting. Through our comprehensive analysis, we identified significant challenges and opportunities in the realm of time-series forecasting, particularly highlighting the need for models that can effectively address both point and probabilistic forecasting across diverse horizons.
Limitations
While our study represents a significant step forward in understanding and evaluating time-series forecasting models, it does come with many limitations. A predominant focus of our work is on empirical analysis, relying heavily on intuitions and experimental observations, which may lack the depth that theoretical foundations could provide. Additionally, our exploration, though extensive, might not encompass all the nuanced factors that influence model performance. By concentrating on major methodological decisions such as AR versus NAR decoding schemes, we may inadvertently overlook other critical aspects that could play a decisive role in forecasting accuracy. Moreover, the datasets employed for evaluation, despite their diversity and relevance to current research threads, may not fully capture the vast spectrum of real-world forecasting challenges. This limitation is particularly pronounced when comparing different foundation models, as their pre-training often involves an even broader array of data, potentially skewing the comparative analysis.
Future Directions
The insights derived from our study open the door to numerous promising research directions. Addressing the shortcomings of AR and NAR decoding schemes, especially in their application across varying forecasting horizons, emerges as a critical area for future exploration. Innovating effective architecture designs that can navigate the intricacies of short-term forecasting challenges and devising efficient methods for long-term probabilistic forecasting stand out as urgent needs. For those time-series foundation models, developing mechanisms to combat error accumulation in AR-based methods could be very helpful. Equally important is the enhancement of models’ abilities to characterize complex data distributions, which could significantly improve the adaptability and effectiveness of foundation models. Beyond these technical endeavors, expanding the scope of datasets used for evaluation to encompass a wider range of real-world scenarios will be crucial for validating the robustness and versatility of future forecasting models. Lastly, integrating theoretical insights with empirical findings could provide a more holistic understanding of model behaviors, contributing to the development of more sophisticated and nuanced forecasting solutions.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Sections 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We release the ProbTS toolkit, documentation, and running scripts at https://github.com/microsoft/ProbTS. The repository includes parameter configurations for benchmarking experiments, ensuring reproducibility of all results presented in the paper.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] All datasets in this paper are split according to standard practices in previous works. We have specified the hyperparameters in Appendix B.3 The detailed configuration file for each baseline is available on our GitHub page.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] We specify a fixed set of random seeds for each experiment to indicate the error bars. We show the mean and standard deviation of evaluation scores collected from multiple runs.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix B.3.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix B.4.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The data utilized in this study is open-source and does not necessitate consent. Acquisition details are available on our GitHub page.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] The data we are using contains no personally identifiable information or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Additional Related Work
A.1 A Comparison with Traditional Models on Covering Essential Forecasting Needs and methodological decisions
@@ Appendix A Additional Related Work
Method Pred. Horizon Paradigm Arch. Design Dec. Scheme
Short Long Point Distr. General Customized AR Non-AR
N-BEATS [52]
Autoformer [70]
Informer [74]
Pyraformer [38]
N-HiTS [12]
LTSF-Linear [72]
PatchTST [48]
TimesNet [69]
iTransformer [39]
DeepAR [58]
GP-copula [57]
LSTM NVP [56]
LSTM MAF [56]
Trans MAF [56]
TimeGrad [55]
CSDI [60]
SPD [8]
TSDiff [33]
This Study
A.2 A Comparison of Pre-trained Time-series Foundation Models
We have incorporated eight recently emerged time series foundation models, namely Lag-Llama [54], Chronos [3], TimesFM [16], Timer [40], MOIRAI [68], UniTS [24], ForecastPFN [19], and TTM [21], into our framework. These foundation models are categorized based on their capabilities, such as zero-shot forecasting, adaptability to varying prediction lengths, and support for probabilistic predictions, as well as their architectural designs, including whether they are auto-regressive and the nature of their backbone networks. Additionally, we have detailed their training processes, including the lengths of prediction horizons used during pre-training and the sizes of look-back windows. These details are summarized in Table 3.
Furthermore, we have compiled a summary of these foundation models’ pre-training and evaluation on several classical time series forecasting datasets. This compilation is presented in Table 4.
Table 3: Foundation Models for Time Series. Zero-shot indicates whether the original work tests zero-shot capabilities. Any-horizon indicates if the same pre-trained model can adapt to prediction tasks of varying lengths. AR denotes if the model performs auto-regressive forecasting. Prob. indicates if the model natively supports probabilistic forecasting. Arch. denotes the model’s backbone architecture: D-O for decoder-only transformer, E-O for encoder-only transformer, E-D for encoder-decoder transformer, and unique for specially designed backbones. Multi-variate indicates if the model explicitly handles multivariate relationships. Pre-train Horizons specifies the forecasting task horizons during pre-training. Look-back Window specifies the context history length settings used in the original experiments.
Model Zero-shot Any-horizon AR Prob. Arch. Multi-variate Pre-train Horizons Look-back Window
Lag-Llama D-O 24$\sim$60 32$\sim$1024
Chronos E-D 64 512
TimesFM D-O -- 512
Timer D-O up to 1440 672
[…]
UniTS E-O -- 60$\sim$720
ForecastPFN E-O 0 50 50 500
TTM Unique 96 512
[…]
Model Solar Wikipedia ETTm1 ETTm2 ETTh1 ETTh2 Electricity Traffic Weather Exchange ILI
[…]
Timer few few few few few few few
A.3 A Comparison with Existing Libraries on the Coverage of Data, Model, and Metrics
ProbTS is a research toolkit designed to advance forecasting research across varied horizons, focusing on both point and distributional forecasting. To achieve these objectives, ProbTS includes state-of-the-art (SOTA) models, comprehensive evaluation protocols (point vs. distributional), and explores different methodological aspects of forecasting models, particularly in terms of distributional estimation methods and decoding schemes (AR vs. NAR). In Table 5, we provide a comprehensive comparison of ProbTS with existing libraries in terms of toolkit functionalities and the benchmarking aspects we aim to investigate.
Table 5: Comparison of Various Time Series Tools.
Tools Features Benchmarking
SOTA Model Dist. Evaluation Short-term Long-term AR vs. NAR Point vs Prob.
Darts [28]
Merlion [7]
Kats [32]
pytorch-forecasting [62]
NeuralForecast [50]
pytorch-transformer-ts [63]
Prophet [61]
sktime [41]
tsai [51]
TSlib [69]
GluonTS [2]
ProbTS
Appendix B More Details on ProbTS
B.1 Data
The data module unifies varied data scenarios to facilitate thorough evaluation and implements standardized pre-processing techniques to ensure fair comparison.
Moreover, we utilize a quantitative approach to visually delineate datasets’ intrinsic characteristics, which employs decomposition to assess trends and seasonality in a time series and evaluate the similarity between data distribution and a Gaussian to depict the complexity of data distribution.
B.1.1 Time-series Forecasting Datasets
@@ B.1.2 Data Visualization
It’s important to note that these visualizations are not selectively chosen or "cherry-picked". We have depicted multiple time-series segments from various time steps, and the observed patterns remain consistent across these different instances.
@@
B.1.3 Quantifying Trend and Seasonality Strengths
Based on the intuition obtained from data visualization, we would like to quantify the strengths of trend and seasonality for a time-series segment with a predefined window size (corresponding to the prediction horizon). Then we can quantify the trend and seasonality strengths at the dataset level by averaging over all time-series segments of a dataset.
To quantify the strengths of trend and seasonality for a fixed-length time-series segment, we draw upon methodologies outlined in the work of [66]. In particular, we employed a time series decomposition model expressed as:
@@ B.1.3 Quantifying Trend and Seasonality Strengths
B.1.4 Quantifying Data Distribution Complexity
@@
Hence, we use the Jensen–Shannon divergence [49] to measure the similarity between the actual value distribution of a time-series segment and a Gaussian distribution fitted to the observed values. Short-term datasets used a window size of 30, while long-term datasets used a size of 336. By averaging the calculated divergence values across all time-series segments of a dataset, we obtain a dataset-level measure of non-Gaussianity. A larger divergence value indicates a larger deviation from a Gaussian distribution in the data.
B.2 Metrics
@@
B.2.1 Metrics for Point Forecasts
@@
$\textrm{MAE}=\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-\hat{x}^{k}_{t}|,$
where $K$ is the number of variates, $T$ is the length of series, $x^{k}_{t}$ and $\hat{x}^{k}_{t}$ denotes the ground-truth value and the predicted value, respectively. For multivariate time series, we also provide the aggregated version:
@@
where $x^{\textrm{sum}}_{t}$ and $\hat{x}^{\textrm{sum}}_{t}$ are the summation across the dimension $K$ of $x^{k}_{t}$ and $\hat{x}^{k}_{t}$, respectively.
@@
$\textrm{NMAE}=\frac{\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-\hat{x}^{k}_{t}|}{\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}|}.$
@@
$\textrm{NMAE}_{\textrm{sum}}=\frac{\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t}|}{\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}|}.$
@@
$\textrm{MSE}=\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}(x^{k}_{t}-\hat{x}^{k}_{t})^{2}.$
@@
$\textrm{MSE}_{\textrm{sum}}=\frac{1}{T}\sum^{T}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}.$
@@
$\textrm{NRMSE}=\frac{\sqrt{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}(x^{k}_{t}-\hat{x}^{k}_{t})^{2}}}{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}|}.$
@@
$\textrm{NRMSE}_{\textrm{sum}}=\frac{\sqrt{\frac{1}{T}\sum^{T}_{t=1}(x^{\textrm{sum}}_{t}-\hat{x}^{\textrm{sum}}_{t})^{2}}}{\frac{1}{T}\sum^{T}_{t=1}|x^{\textrm{sum}}_{t}|}.$
Mean Absolute Scaled Error (MASE)
The Mean Absolute Scaled Error (MASE) divideds the MAE of forecasted values by MAE of the in-sample one-step naive forecast, which is a scale-invariant metrics:
$\textrm{MASE}=\frac{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-\hat{x}^{k}_{t}|}{\frac{1}{K\times T}\sum^{K}_{k=1}\sum^{T}_{t=1}|x^{k}_{t}-x^{k}_{t-1}|}.$
B.2.2 Metrics for Distributional Forecasts
@@
The Continuous Ranked Probability Score (CRPS) [46] quantifies the agreement between a cumulative distribution function (CDF) $F$ and an observation $x$, represented as:
@@
where $\mathds{I}\{x\leq z\}$ denotes the indicator function, equating to one if $x\leq z$ and zero otherwise.
@@
B.3 Baselines
To ensure the integrity of the results, ProbTS adheres to a standard implementation process, employing unified data splitting, standardization techniques, and adopting fair settings for hyperparameter tuning across all methods.
Implementation Details
ProbTS was developed using PyTorch Lightning [23]. During training, we sampled 100 batches per epoch and limited training to 50 epochs, using the CRPS metric for checkpointing. All experiments employed the Adam optimizer and were run on single NVIDIA Tesla V100 GPUs with CUDA 11.3. To enable evaluation of distribution-level metrics, we conducted 100 samplings to calculate metrics on the test set.
Following the most commonly adopted settings [72, 48, 69], in the long-term forecasting context, all of the models are following the same experimental setup with prediction length $T\in\{24,36,48,60\}$ for ILI-L dataset and $T\in\{96,192,336,720\}$ for other datasets. Note that the lookback window here is 96 for all the models, to ensure a fair comparison. In the short-term forecasting context, the length of the lookback window is the same as the forecasting horizons, which are 30 for Exchange-S dataset and Wikipedia-S dataset, and 24 for the rest, the same as [57].